### v2

# Polynomial parametrization of some interesting knotted surfaces

Abstract.
We discuss methods to construct a polynomial parametrization of some interesting knotted surfaces (knotted spheres, knotted tori and knotted planes) and provide examples.
Key words and phrases:
Spun knots, twist spun knots, welded knots, ribbon torus knots, polynomial parametrization
2020 Mathematics Subject Classification:
$57K10,57K12,57K14,57K45$
1. Introduction
Knots have a rich and long history and have been used for various activities like fishing, hunting, and building. The knots that we see are one dimensional objects inside our three dimensional space. To ensure a most simple form of a knot, one simplifies it without cutting it and mathematicians call it ambient isotopy [5]. Many topological and geometric properties of a knot can be inferred if one can produce a parametrization of the knot using simple functions. Knot parametrizations are studied for classical knots by many mathematicians [18, 9, 14, 6]. Similar to a rope getting knotted, even surfaces (such as plane, sphere, torus) can get knotted. Surface knots are seen in computer graphics, physics and material science. In material science and textile, microscale knots [19] are example of surface knots. It is believed that these knotted surfaces provide stability and durability to the structure [15]. Thus interesting knotted surfaces may help material scientists design different and aesthetically elegant structures. Knotted surfaces are difficult to visualize due to their complex topology. Thus if one can represent these knotted surfaces by nice parametric equations, one is able to judge its geometric properties that in turn provides an information on the strength it can create to the material it is used for.
In this paper, we have provided examples of knotted spheres, knotted tori and knotted planes. Our examples rely on constructing these surfaces with the help of one dimensional knots such as classical knots, welded knots and long knots. We provide mathematical formulation to capture these constructions and use the knot parametrization of classical and long knots to parametrize these knotted surfaces. Thus, we can provide plenty of examples to be utilized by the scientists and engineers for their practical purposes.
This paper is organized as follows. Section 2 includes the basic definitions required to study surface knots. Section 3 is divided into two subsections, Section 3.1 and Section 3.2 defines two constructions, spinning and twist spinning respectively for obtaining 2-knots from classical knots and discuss their parametrizations. In Section 4, we provide a brief exposition of the welded knots and its relationship with ribbon torus knots and then discuss the method to find a polynomial parametrization of the ribbon torus knots using that relationship. In Section 5, we discuss methods of constructing knotted planes from long classical knots. For each construction we have included the images generated from Mathematica.
2. Basics on surface knots
To study surface knots rigorously, we first recall several foundational concepts related to embeddings of manifolds. These definitions are essential for understanding the structure and classification of surfaces embedded in four-dimensional space.
Definition 2.1.
Let $N$ be an $n$-dimensional manifold embedded in an $m$-dimensional manifold $M$. We call an embedding $f:N\rightarrow M$ to be proper if every compact subset $K\subset f(N)$ has a compact preimage $f^{-1}(K)=\{x\in N|f(x)\in K\}$. In other words, the manifold $f(N)$ is proper if $\partial f(N)=f(N)\cap\partial M$.
To ensure that surface embeddings behave nicely in small neighbourhoods, we need a notion of local flatness.
Definition 2.2.
Let $f:N\to M$ be a proper embedding of smooth manifolds. This embedding is said to be locally flat at a point $x\in M$ if there exists a regular neighborhood $U$ of $x$ in $M$ such that the pair $(U,U\cap f(N))$ is homeomorphic to the standard ball pair $(D^{m},D^{n})$, where $D^{k}$ is the unit ball in $\mathbb{R}^{k}$, centred at the origin. We say that $f$ is locally flat if it is locally flat at every point of $M$.
Not all proper embeddings are locally flat. Examples of non-locally flat embeddings can be found in [11], highlighting the necessity of this condition in knot theory.
With these foundational notions established, we now introduce the definitions of surface knots and surface links.
Definition 2.3.
A surface link is a proper, locally flat embedding of a closed surface $F$ in $\mathbb{R}^{4}$ or $S^{4}$. If $F$ is a connected surface then it is called a surface knot. When $F=S^{2}$, it is called a 2-knot or $2$-dimensional knot.
Surface knots can also appear in non-compact forms, leading to the notion of long surface knots.
Definition 2.4.
A long $2$-knot is a proper, locally flat embedding of $\mathbb{R}^{2}$ into $\mathbb{R}^{4}$ that is standard outside a compact region.
Remark 2.1.
Every $2$-knot in $S^{4}$ naturally gives rise to a long $2$-knot in $\mathbb{R}^{4}$ by removing a point from the embedded $S^{2}$ and its image from $S^{4}$. This converts the compact embedding into a non-compact one, effectively yielding a long version of the $2$-knot.
Now, we recall an interesting class of knotted surfaces, called ribbon surface knots.
Definition 2.5.
Let $M$ be an compact $3$-manifold immersed in $\mathbb{R}^{4}$ and $f:M\rightarrow\mathbb{R}^{4}$ be the immersion map. A connected component $D$ of the double point set is called a ribbon singularity if the preimage $f^{-1}(D)$ is the union of $2$-disks $D_{1}$ and $D_{2}$ such that $D_{1}$ is a properly embedded 2-disk in $M$ and $D_{2}$ is an embedded $2$-disk in the interior of $M$.
Definition 2.6.
Let $\Delta=\Delta_{1}\sqcup\cdots\sqcup\Delta_{k}$ and $\mathcal{B}=\mathcal{B}_{1}\sqcup\cdots\sqcup\mathcal{B}_{l}$ be two disjoint collections of $3$-balls embedded in $\mathbb{R}^{4}$, with each component of $\Delta$ and $\mathcal{B}$ is called a base and a band respectively. We parametrize each band by $h_{i}:D^{2}\times[0,1]\rightarrow\mathbb{R}^{4}$ such that
  1. (1)
    For each $i=1,2,\cdots,l$,
    $\partial\Delta\cap h_{i}(D^{2}\times[0,1])=h_{i}(D^{2}\times\partial[0,1]),\quad\text{and}$
  2. (2)
    If $h_{i}(D^{2}\times[0,1])$ intersects a base, then it intersects this base in meridional 2-disks that are prpoerly embedded in the band and contained in the interior of the base:
    $h_{i}(D^{2}\times[0,1])\cap\Delta_{j}=h_{i}(D^{2}\times\{t\})$
    for $t\in(0,1)$. And all these intersections are ribbon singularities (Fig. 1).
Now the surface given by $F=\left[\partial\Delta\,\setminus\,\cup_{i=1}^{l}h_{i}(D^{2}\times[0,1])\,\right]\bigcup\,\cup_{i=1}^{l}h_{i}(\partial D^{2}\times[0,1]),$ is called a ribbon surface knot (Fig. 2).
  • If l=k1l=k-1 then $F$ is homeomorphic to a $2$-sphere and $F$ is called a ribbon 2-knot or a ribbon sphere.
  • If $F$ is homeomorphic to a torus then we call it a ribbon torus-knot.
  • The pair $(\Delta,\mathcal{B})$ is called a ribbon presentation of $F$ with $l$ fusion bands.
The intersections of type (1) and (2) are depicted in Fig. 1 for classical case (bottom) and surface case (top). Fig. 2 shows a broken surface diagram of a ribbon surface knot where parts of the surface is removed to indicate the over/under information similar to classical knot theory.
Figure 1. Ribbon presentation in dimension one and two
Figure 2. A broken surface diagram of a ribbon surface knot
For a detailed exposition of surface knots and their topological subtleties, we refer the reader to [4, 8, 11].
In the upcoming sections we will discuss knotted spheres, knotted tori and knotted planes. Our main focus is to provide parametric equations, so in each case, we have attempted only those classes of knotted surfaces which arise naturally from the one dimensional knots so that their parametrization can be used.
3. Parametrizing knotted spheres
A two dimensional sphere can get knotted inside a four dimensional space $\mathbb{R}^{4}$ or $S^{4}$. One can study different isotopy classes with the help of invariants. These $2$-dimensional knots are also known as $2$-knots. Since these knots are in $\mathbb{R}^{4}$ it is difficult to visualize them. Moreover, there is no uniform method to enumerate them and study them in order. Simplest methods to visualize a knotted $S^{2}$ is through it projection into some three dimensional space or to look at the movie [4]. Our goal is to produce parametric equations that represent interesting knotted spheres. It is known that certain operations (spinning, twist spinning, roll spinning [7]) performed on a classical knot inside $\mathbb{R}^{3}$ result into a knotted sphere inside $\mathbb{R}^{4}$. Thus with a parametric equations of a classical knot in hand, if we can formulate these operations mathematically, we obtain parametrization of such knotted spheres. We have succeeded in constructing spinning and $d$-twist ($d\geq 0$) spinning operations and the knots obtained are called spun $2$-knots and $d$-twist spun $2$-knots. We describe these constructions in Section 3.1 and Section 3.2 respectively.
3.1. Spun $2$-knots
E. Artin [2] introduced a way to construct $2$-knots by spinning a knotted arc embedded in the half-space $\mathbb{R}^{3}_{+}$ around $\mathbb{R}^{2}$. The resulting $2$-knots are called spun 2-knots. Spun knots represent the simplest class of $2$-knots. Their construction is described as follows. In $\mathbb{R}^{4}$, consider the upper half space $\mathbb{R}^{3}_{+}=\left\{(x,y,z,0):z\geq 0\right\}$ with the boundary $\partial\mathbb{R}^{3}_{+}=\mathbb{R}^{2}=\left\{(x,y,0,0)\right\}.$ Now, locus of a point $x=(x,y,z,0)$ in $\mathbb{R}^{3}_{+}$, rotated in 4-space about $\mathbb{R}^{2}$ is given by
$\left\{(x,y,z\,\cos\theta,z\,\sin\theta)\,\middle|\,0\leq\theta\leq 2\pi\right\}.$
Figure 3. Spinning
To construct a spun 2-knot, we choose a properly, locally flatly embedded arc $k$ in $\mathbb{R}^{3}_{+}$ i.e $k$ is embedded in $\mathbb{R}^{3}_{+}$ locally flatly and intersects $\partial\mathbb{R}^{3}_{+}$ transversely only at the endpoints (Fig. 3). Then we spin $\mathbb{R}^{3}_{+}$ along $\mathbb{R}^{2}$ by $360^{o}$ and the continuous locus of the arc $k$ produces the spun 2-knot given by
$\left\{(x,y,z\,\cos\theta,z\,\sin\theta)\,\middle|\,(x,y,z)\in k,0\leq\theta\leq 2\pi\right\}.$
Theorem 3.1.
Given a classical knot $K$, there exist polynomials $f(s,t)$, $g(s,t),h(s,t)$ and $p(s,t)$ in two variables $s$ and $t$ such that for some interval $[a,b]$ the image of $a,b]\times[0,2\pi]$ under the map $\phi:\mathbb{R}^{2}\to\mathbb{R}^{4}$ defined by
$\phi((s,t))=\left(f(s,t),g(s,t),h(s,t),p(s,t)\right)$
is isotopic to the spun of $K$.
Proof.
Let us choose a polynomial representation $\phi$ of $K$ such that $\phi:[a,b]\rightarrow\mathbb{R}^{3}_{+}$ of the knotted arc $k$ is given by $\phi(t)\rightarrow\left(f(t),g(t),h(t)\right)$ such that $h(a)=0=h(b)$ and $h(t)>0$ for $t\in(a,b)$. Now the spinning construction gives a map $F:[a,b]\times[0,2\pi]\to\mathbb{R}^{4}$ defined by
$F(s,t)=\left(f(t),\,g(t),\,\,h(t)\,\cos s,\,h(t)\,\sin s\right)$
that represents $spun(K)$. Using a point set topology argument, one can prove that the image of $F$ in $\mathbb{R}^{4}$ is homeomorphic to $S^{2}$. Now, we replace the trigonometric functions $\cos s$ and $\sin s$ with their Chebyshev approximations inside the interval $0,2\pi] [$1]. Let the Chebyshev approximations of $\cos s$ and $\sin s$ inside the interval $0,2\pi]$ be denoted by $C(s)$ and $S(s)$, respectively. Then by choosing $f(s,t)=f(t)$, $g(s,t)=g(t)$, h(s,t)=h(t)C(s)h(s,t)=h(t)\,C(s) and p(s,t)=h(t)S(s)p(s,t)=h(t)\,S(s), we obtain a polynomial parametrization of $spun(K)$. This completes the proof. ∎
Example 3.1 (The spun trefoil knot).
Let us take the following polynomial representation of the long trefoil knot [18] given by
$\left\{\left(t^{3}-3t,t^{5}-10t,t^{4}-4t^{2}\right)\,\middle|\,t\in\mathbb{R}\right\}.$
We change the function $(t^{4}-4t^{2})$ to $(-t^{4}+4t^{2}+3)$ which has real roots t=2.1554t=-2.1554 and $t=2.1554$. Therefore, the knotted arc $k$ (Fig. 4) is given by
$\left\{\left(t^{3}-3t,t^{5}-10t,-t^{4}+4t^{2}+3\right)\,\middle|\,t\in[-2.1554,2.1544]\right\}.$
and the spun trefoil knot is given by
$\left\{\left(t^{3}-3t,t^{5}-10t,(-t^{4}+4t^{2}+3)\,C(s),(-t^{4}+4t^{2}+3)\,S(s)\right)\right\}$
for $t\in[-2.1554,2.1544],s\in[0,2\pi]$. For the final polynomial parametrization we take the Chebyshev approximations $C(s)$ and $S(s)$ of $\cos s$ and $\sin s$ in $0,2\pi]$ respectively as follows:
[…]
Mathematica plot for a projection of the spun trefoil are shown in Fig. 5(a). In Fig. 5(b), an inside view is also given where it can be seen that the knotted arc does not deform while spinning around $\mathbb{R}^{2}$.
Figure 4. Knotted arc of the long trefoil knot
(a) Projection on a hyperplane
(b) An inside view
Figure 5. The spun trefoil knot
Example 3.2 (The spun figure-eight knot).
Using the parametrization for the long figure eight knot given in [3] (Fig. 6) and following the same procedure as above, we obtain a polynomial parametrization of the spun figure eight knot given by
$\left\{\left(f(t),g(t),h(t)\,C(s),h(t)\,S(s)\,\right)\,\middle|\,t\in[-3.7934,3.7934],s\in[0,2\pi]\right\}$
$\text{ where}\qquad f(t)$ $=\frac{2}{5}\left(t^{2}-7\right)\left(t^{2}-10\right)t,$
$g(t)$ $=\frac{1}{10}t\left(t^{2}-4\right)\left(t^{2}-9\right)\left(t^{2}-12\right),$
$h(t)$ $=(20-13t^{2}-t^{4}).$
and $C(s)$, $S(s)$ are the Chebyshev approximations of $\cos s$ (Equation 1) and $\sin s$ (Equation 2) respectively. Fig. 7 shows the Mathematica plot of a projection of the spun figure eight knot on XZW-plane.
Figure 6. Knotted arc of long figure eight knot
Figure 7. The spun figure eight knot
Remark 3.2.
Thus for every classical knot $K$, we will be able to construct a spun $2$-knot parametrized by polynomial functions.
3.2. $d$-twist spun $2$-knots
E. C. Zeeman generalized Artin’s spinning construction to twist spinning in 1965 [21]. In this case, we imagine the knotted part of $k$ inside a 3-ball as in Fig. 8. Now, to include twisting in the spinning construction we rotate the ball $d$ times around its own axis while rotating $\mathbb{R}^{3}_{+}$ around $\mathbb{R}^{2}$ once. Position of the knotted arc after $d$ twists should match its initial position after completing the rotation around $\mathbb{R}^{2}$. This way we get a 2-sphere in $\mathbb{R}^{4}$, called $d$-twist spun 2-knot. By definition, for $d=0$ we get a spun knot.
Figure 8. Twist spinning
Theorem 3.3.
Given a classical knot $K$, there exist polynomials $f(t,\theta)$, $g(t,\theta),h(t,\theta)$ and $p(t,\theta)$ in two variables $t$ and $\theta$ such that for some interval $[a,b]$, the image of $a,b]\times[0,2\pi]$ under the map $\phi:\mathbb{R}^{2}\to\mathbb{R}^{4}$ defined by
$\phi(t,\theta)=\left(\bar{f}(t,\theta),\bar{g}(t,\theta),\bar{h}(t,\theta),\bar{p}(t,\theta)\right)$
is isotopic to the $d$-twist spun of $K$.
Proof.
We start with the parametrization $\phi:[a,b]\rightarrow\mathbb{R}^{3}_{+}$ of the knotted arc $k$, given by
$\phi(t)\rightarrow\left(f(t),g(t),h(t)\right)$
where $h(a)=0=h(b)$ and $h(t)>0$ for $t\in(a,b)$. The endpoints of $k$ on the boundary $\mathbb{R}^{2}$ is given by $(f(a),g(a),0)$ and $(f(b),g(b),0)$. Now we find an interval $a^{\prime},b^{\prime}]\subset[a,b]$ where all the crossings of $K$ lie.
In the twist spinning construction, the knotted part of the arc is contained in a 3-ball in $\mathbb{R}^{3}_{+}$ and is rotated around the axis of the ball. To achieve this first we choose the axis of rotation as a line segment $PQ$ parallel to the $XY$ plane joining two points on the knotted arc, say $P:=(f(t_{1}),g(t_{1}),h(t_{1}))$ and $Q:=(f(t_{2}),g(t_{2}),h(t_{2}))$, where $a^{\prime},b^{\prime}]\subset[t_{1},t_{2}]$ and $h(t_{1})=h(t_{2})=c$ (Fig. 9). The value of $c$ is chosen so that the knotted part of the arc does not intersect the $XY$ plane while rotating around $PQ$. By Rodrigues’ rotation formula [16], the rotation around $PQ$ is represented by the following matrix.
$\mathbf{R}^{\prime}=\mathbf{T_{c}*R*T_{c}^{-1}}=\mathbf{T_{c}*R*T_{-c}},$
where $\mathbf{T_{c}}$ is the translation matrix along Z axis which will send $(x,y,z)$ to $(x,y,z+c)$ and $\mathbf{R}$ gives the rotation around the line $P^{\prime}Q^{\prime}$ on $XY$ plane, parallel to $PQ$, joining the points $P^{\prime}=(f(t_{1}),g(t_{1}),0)$ and $Q^{\prime}=(f(t_{2}),g(t_{2}),0)$ on the knotted arc.
Figure 9. Rotating about PQPQ
The rotation matrix $\mathbf{R}$ is defined as follows: If $\mathbf{v}$ is a vector in $\mathbb{R}^{3}$ and $\mathbf{k}$ is a unit vector describing an axis of rotation around which $\mathbf{v}$ rotates by an angle $\phi$ according to the right-hand rule, the Rodrigues formula for the rotated vector $\mathbf{v_{rot}}$ is given by
$\mathbf{v}_{rot}=\mathbf{v}\,\cos\phi+(\mathbf{k}\times\mathbf{v})\sin\phi+\mathbf{k}\cdot(\mathbf{k}\cdot\mathbf{v})(1-\cos\phi).$
In this case, $\mathbf{k}$ is along the line joining $(f(t_{1}),g(t_{1}),0)$ and $(f(t_{2}),g(t_{2}),0)$. So,
$\mathbf{k}=\Big(\frac{f(t_{2})-f(t_{1})}{N},\frac{g(t_{2})-g(t_{1})}{N},0\Big),$
where $N=\sqrt{(f(t_{2})-f(t_{1}))^{2}+(g(t_{2})-g(t_{1}))^{2}}.$
For simplification, let us denote
$f_{21}:=f(t_{2})-f(t_{1})$
$g_{21}:=g(t_{2})-g(t_{1}).$
Then the rotation matrix through an angle $\phi$ counterclockwise around the axis $\mathbf{k}$ is
$\mathbf{R=I+\sin\phi\;K+(1-\cos\phi)\;K^{2}},$
where
$\mathbf{K}=\begin{pmatrix}0&-k_{z}&k_{y}\\ k_{z}&0&-k_{x}\\ -k_{y}&k_{x}&0\end{pmatrix}=\begin{pmatrix}0&0&\dfrac{g_{21}}{N}\\[14.22636pt] 0&0&-\dfrac{f_{21}}{N}\\[14.22636pt] -\dfrac{g_{21}}{N}&\dfrac{f_{21}}{N}&0\end{pmatrix}$.
and the rotation matrix around $PQ$ is given by,
$\mathbf{R}^{\prime}=\mathbf{T_{c}*R*T_{c}^{-1}}=\mathbf{T_{c}*R*T_{-c}},$
where
$\mathbf{R}=\begin{pmatrix}\dfrac{f_{21}^{2}\;+\;g_{21}^{2}\;\cos\phi}{N^{2}}&\dfrac{f_{21}\;g_{21}\;(1-\cos\phi)}{N^{2}}&\dfrac{g_{21}\;\sin\phi}{N}&0\\[14.22636pt] \dfrac{f_{21}\;g_{21}\;(1-\cos\phi)}{N^{2}}&\dfrac{f_{21}^{2}\;\cos\phi\;+\;g_{21}^{2}}{N^{2}}&-\dfrac{f_{21}\;\sin\phi}{N}&0\\[14.22636pt] -\dfrac{g_{21}\;\sin\phi}{N}&\dfrac{f_{21}\;\sin\phi}{N}&\cos\phi&0\\[14.22636pt] 0&0&0&1\end{pmatrix}$
and $\mathbf{T_{c}}$ is the translation matrix along Z-axis which sends $(x,y,z)$ to $(x,y,z+c)$, given by
$T_{c}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&c\\ 0&0&0&1\end{pmatrix}$.
Then,
$\mathbf{R}^{\prime}=\begin{pmatrix}\dfrac{f_{21}^{2}\;+\;g_{21}^{2}\;\cos\phi}{N^{2}}&\dfrac{f_{21}\;g_{21}\;(1-\cos\phi)}{N^{2}}&\dfrac{g_{21}\;\sin\phi}{N}&-\dfrac{c\;g_{21}\;\sin\phi}{N}\\[14.22636pt] \dfrac{f_{21}\;g_{21}\;(1-\cos\phi)}{N^{2}}&\dfrac{f_{21}^{2}\;\cos\phi\;+\;g_{21}^{2}}{N^{2}}&-\dfrac{f_{21}\;\sin\phi}{N}&\dfrac{c\;f_{21}\;\sin\phi}{N}\\[14.22636pt] -\dfrac{g_{21}\;\sin\phi}{N}&\dfrac{f_{21}\;\sin\phi}{N}&\cos\phi&-c\;\cos\phi+c\\[14.22636pt] 0&0&0&1\end{pmatrix}$.
Thus, the rotation around $PQ$ is given by the parametrization :
$(t,\phi)\longrightarrow\left\{\left(f^{\prime}(t,\phi),g^{\prime}(t,\phi),h^{\prime}(t,\phi)\right)\,\middle|\,\begin{aligned} a\leq&t\leq b\\ 0\leq&\phi<2\pi\end{aligned}\,\right\}$ (3)
where
$f^{\prime}(t,\phi)$ $=\frac{\left(f_{21}^{2}\;+\;g_{21}^{2}\;\cos\phi\right)\,\mathbf{f(t)}+f_{21}\;g_{21}\;(1-\cos\phi)\,\mathbf{g(t)}+N\;g_{21}\,\sin\phi\,(\mathbf{h(t)}-c)}{N^{2}}$ (4)
$g^{\prime}(t,\phi)$ $=\frac{f_{21}\;g_{21}\;(1-\cos\phi)\,\mathbf{f(t)}+\left(f_{21}^{2}\;\cos\phi\;+\;g_{21}^{2}\right)\,\mathbf{g(t)}+N\;g_{21}\,\sin\phi\,(c-\mathbf{h(t)})}{N^{2}}$ (5)
$h^{\prime}(t,\phi)$ $=\frac{-g_{21}\;\sin\phi\;\mathbf{f(t)}+f_{21}\sin\phi\;\mathbf{g(t)}+N\,\cos\phi\,\mathbf{h(t)}+N\,c\,(1-\cos\phi)}{N}$ (6)
Now, in the twist-spinning construction, there is no rotation outside the $3$-ball that contains the knotted part. This is same as not rotating the knotted arc outside the endpoints of $PQ$. We achieve this by combining the rotation functions with a bump function.
We have, $a,b]=[a,t_{1}]\cup[t_{1},a^{\prime}]\cup[a^{\prime},b^{\prime}]\cup[b^{\prime},t_{2}]\cup[t_{2},b]$ (Fig. 9), where
  1. (1)
    $a^{\prime},b^{\prime}]$ contains all the crossings of the knot.
  2. (2)
    $t_{1},t_{2}$ corresponds to the endpoints of $P$ and $Q$. So, $a,t_{1}]$ and $t_{2},b]$ correspond to the parts of the arc outside the axis $PQ$.
Figure 10. Bump function B(t).
Therefore, we define a bump function which takes the value $1$ in $a^{\prime},b^{\prime}]$ and 0 outside $t_{1},t_{2}]$. First, consider the function
$F(t)=\begin{cases}e^{-\frac{1}{t}}\,,&t>0\\ 0\,,&t\leq 0.\end{cases}$
Then define the bump function, $B(t)$ as follows:
$B(t)=\frac{F(d_{1}-t^{2})}{F(t^{2}-d_{2})+F(d_{1}-t^{2})},$
where $d_{1},d_{2}\in\mathbb{R}^{+}$ are chosen in such a way that
$B(t)=\begin{cases}1\,,&t\in[a^{\prime},b^{\prime}]\\ 0\,,&t\in[a,t_{1}]\cup[t_{2},b]\\ \in(0,1)\,,&\text{otherwise}.\end{cases}.$
See Fig. 10. Then, for $t\in[a,b]$, we define
$\tilde{f}(t,\phi)=$ $\,B(t)\,f^{\prime}(t,\phi)+(1-B(t))\,f(t),$
$\tilde{g}(t,\phi)=$ $\,B(t)\,g^{\prime}(t,\phi)+(1-B(t))\,g(t),$
$\tilde{h}(t,\phi)=$ $\,B(t)\,h^{\prime}(t,\phi)+(1-B(t))\,h(t).$
Now, the knotted arc is rotating $d$ times around the $PQ$ axis in $\mathbb{R}^{3}_{+}$ while rotating around the $XY$ plane in $\mathbb{R}^{4}$. Therefore, the rotation angle around $PQ$ is $d$ times the rotation angle around the $XY$ plane i.e. $\phi=d\,\theta.$
Hence, the parametrization for $d$-twist spun knot is given by
$(t,\theta)\longrightarrow\left\{\left(\tilde{f}(t,d\theta)\,,\,\tilde{g}(t,d\theta)\,,\,\tilde{h}(t,d\theta)\,\cos\,\theta\,,\,\tilde{h}(t,d\theta)\,\sin\,\theta\right)\,\middle|\,\begin{aligned} a\leq&t\leq b\\ 0\leq&\theta<2\pi\end{aligned}\right\}.$ (7)
Here, $\tilde{f},\tilde{g},\tilde{h}$ are linear combinations of the functions $f,g,h$, cosine, sine and the bump function $B(t)$. Replacing cosine, sine and $B(t)$ with their corresponding Chebyshev polynomial approximations in $0,2\pi]$, we obtain a polynomial parametrization of $d$-twist spun knots given by
$\left\{\left(\bar{f}(t,\theta)\,,\,\bar{g}(t,\theta)\,,\,\bar{h}(t,\theta)\,,\,\bar{p}(t,\theta)\right)\,\middle|\,a\leq t\leq b,\,0\leq\theta<2\pi\right\}.$ (8)
This completes the proof. ∎
Example 3.3 ($d$-twist spun trefoil).
Start with a polynomial parametrization of the long trefoil knot (Fig. 11), given by
$x(t)$ $=t^{3}-3t$
$y(t)$ $=t^{5}-10t$
$z(t)$ $=-t^{4}+4t^{2}+16.$
Solving for $t_{1},t_{2}$ such that $x(t_{1})=x(t_{2})$ and $y(t_{1})=y(t_{2})$ we can get the crossing data and solving $z(t)=0$ will give the interval for the arc. Therefore, we have
$[a,b]$ $=[-2.54404,2.54404],$
$[a^{\prime},b^{\prime}]$ $=[-1.946,1.946],$
$P$ $=(f[-2.19],g[-2.19],h[-2.19]),$
$Q$ $=(f[2.19],g[2.19],h[2.19]),$
$c$ $=h[2.19].$
Figure 11. knotted trefoil arc with the axis of rotation
Here we define the bump function as
$B(t)=\frac{F(4.8-t^{2})}{F(t^{2}-3.8)+F(4.8-t^{2})}.$
Fig. 12 and 13 shows how the rotation about $PQ$ is restricted using the bump function.
Figure 12. Rotating the long trefoil arc about PQPQ
Figure 13. Restricting rotation using bump function.
Some of the projections of the $d$-twist spun trefoil on a hyperplane for the values $d=0,1,2,5,10,20$ are given below. In Fig. 14(a), it is clear that $d=0$ represents Artin’s spun knot. Fig. 14(b) and Fig. 15(b) illustrate the deformation of the knotted arc while rotating around $\mathbb{R}^{2}$ in the twist spun knot construction, which differs from the Artin’s spun knot construction. The Mathematica notebook for $d$-twist spun knot parametrization can be found in [20].
(a) 0-twist
(b) 0-twist: inside view
Figure 14. $d=0$
(a) 1-twist
(b) 1-twist: inside view
Figure 15. $d=1$
(a) 2-twist
(b) 5-twist
(c) 10-twist
(d) 20-twist
Figure 16. $d=2,5,10,20.$
4. Parametrizing knotted tori
Knotted tori are ambient isotopy classes of embeddings of a torus $S^{1}\times S^{1}$ inside $\mathbb{R}^{4}$. As there can be various classes and it is very difficult to capture them in general. However, a special class of knotted tori known as ribbon torus knots (Definition 2.6) are possible to parametrize due to their connection with the welded (virtual with additional moves) knots. In Section 4.1, we discuss briefly about welded knot theory and in Section 4.2, we provide a method to parametrize a ribbon torus knot.
4.1. Welded knot theory
Welded knot theory is obtained from virtual knot theory, a combinatorial generalization of classical knot theory [10, 17]. Here, a new type of crossing, called virtual crossing, is introduced in the knot diagram(Fig. 17). A knot diagram with classical and virtual crossings is called a virtual knot diagram up to generalized Reidemeister moves.
(a) Classical crossings
(b) Virtual crossing
Figure 17. Two types of crossings
There are two other local moves, called over and under forbidden moves (Fig. 18). Allowing both on a virtual diagram results in a trivial diagram. But allowing only the over forbidden move generates a new non-trivial knot theory, called welded knot theory. And two virtual diagrams, $K$ and $K^{\prime}$, are called welded equivalent, denoted by K $\overset{w}{\sim}K^{\prime}$ if they are related by a finite sequence of generalized Reidemeister moves and the over forbidden move [17]. When we are in welded knot theory, we shall refer to a virtual crossing as a welded crossing. A virtual (welded) knot diagram can be thought to be obtained from a classical knot diagram in which a few crossings are converted as virtual (resp. welded) crossing. We loosely say that these crossings have been virtualized (resp. welded).
Figure 18. Forbidden moves
S. Satoh showed a correspondence between ribbon torus knots and virtual knots by defining a map called Tube map [17]. Let $K$ be a virtual knot diagram on $XY$-plane. We construct a surface diagram corresponding to $K$ in the following way:
(a) Near the classical crossing
(b) Near the virtual crossing
Figure 19. Tube map near the crossings
  • At every point of the diagram, we will put a circle in the $Z$ direction.
  • At a classical crossing, the tube corresponding to the under arc will penetrate through the tube corresponding to the over-arc (Fig. 19(a)).
  • At the virtual crossings, the tubes are free from the over/under information of the diagram in $\mathbb{R}^{3}$, and they can move independently (Fig. 19(b)).
So, Tube(K) is defined as the ribbon torus knot with the surface diagram described above for $K$. He also proved the following.
Proposition 4.1.
[17] $K\overset{w}{\sim}K^{\prime}\implies Tube(K)\cong Tube(K^{\prime}).$
4.2. Parametrizing a ribbon torus knot
As discussed in Section 4.1, every ribbon torus knot is isotopic to tube of some welded knot. Welded knots can be created with the help of classical knots by welding few crossings. Thus from one classical knot one can obtain various non-isotopic welded knots and hence by performing tube operation, we can obtain ribbon torus knots. In Section 4.1, we have seen that at a classical crossing, the tube corresponding to the under-arc will penetrate through the tube corresponding to the over-arc and at a virtual crossing the tubes are free from the over/under information of the diagram in $\mathbb{R}^{3}$, and they can move independently. Below, we describe how to achieve this in parametrization. We prove the following theorem.
Theorem 4.2.
Given a classical knot $K$ with given trigonometric parametrization $\phi^{\prime}:[0,2\pi]\rightarrow\mathbb{R}^{3}$ given by
$\phi^{\prime}(t)=\left(f(t),g(t),h(t)\right),$
there exists functions $S_{c}(t)$ and $S_{w}(t)$ such that the image of $0,2\pi]\times[0,2\pi]$ under the map $\tilde{\Psi}:\mathbb{R}^{2}\to\mathbb{R}^{4}$ defined by
$\tilde{\Psi}(t,s)=\left(f(t),\,(r-d_{1}S_{c}(t))\,g(t)+\cos s,\,(r-d_{1}S_{c}(t))\,g(t)+\sin s+d_{2}S_{w}(t),\,h(t)\right),$
where $r,d_{1},d_{2}\in\mathbb{R}^{+}$, is isotopic to the $Tube(K)$.
Proof.
Take the projection of $K$ on $XY$ plane given by $\{f(t),g(t)\,|\,t\in[0,2\pi]\}$ and let there be $n$ number of classical crossings, denoted by $c_{1},c_{2},\cdots,c_{n}$, and $m$ number of welded crossings, denoted by $w_{1},w_{2},\cdots,w_{m}$. Additionally, let $\{t_{i},s_{i}\}$ be the pair of values of $t$ where $t_{i}$ corresponds to the overcrossing point and $s_{i}$ corresponds to the undercrossing point of $c_{i}$ respectively for $i=1,\cdots,n$. Also, let $\{\bar{t}_{j},\bar{s}_{j}\}$ be the pair of values of $t$ where $\bar{t}_{j}$ corresponds the first encounter of the crossing $w_{j}$ and $\bar{s}_{j}$ corresponds the second encounter of the crossing $w_{j}$ for $j=1,\cdots,m$.
First, putting a circle with a fixed radius $r\in\mathbb{R}_{+}$ at each point of the projection is given by
$\left\{\big(f(t),\,r\,g(t)+\cos s,\,r\,g(t)+\sin s\big)\,\middle|\,t,\,s\in[0,2\pi]\right\}.$
Now, we construct functions for shrinking the radius of the lower tube and displacing the tubes from each other, denoted by $S_{c}(t)$ and $S_{w}(t)$ respectively.
Take intervals of the same length, say $L$, around each $t_{i}$ and $s_{i}$ such that no two intervals intersect. Now, we define the bump function with centre $c$ and width $w$ (Fig. 20) by
$B(x,c,w)=\begin{cases}\exp\left(-\frac{w^{2}}{w^{2}-(x-c)^{2}}\right)&\text{if}\;|x-c|<w\\ 0&\text{otherwise.}\end{cases}$
Figure 20. The bump function $B(x,c,w)$ with center $c=0$ and width $w=\frac{\pi}{2}$
The function $S_{c}(t)$ is defined by summing the functions $B(x,c,\frac{L}{2})$ with centres at the undercrossing points $s_{i}$ of all the classical crossings. On the other hand, the function $S_{w}(t)$ is defined by taking an alternating sum of $B(x,c,\frac{L}{2})$ with centres at the welded crossing points $\bar{t}_{j}$ and $\bar{s}_{j}$ of all welded crossings. Thus,
$S_{c}(x)=$ $\sum_{i=1}^{n}B(x,s_{i},\frac{L}{2}),$
$S_{w}(x)=$ $\sum_{j=1}^{m}\left(B(x,\bar{t}_{j},\frac{L}{2})-B(x,\bar{s}_{j},\frac{L}{2})\right).$
The parametrization of the ribbon torus $Tube(K)$ is given by $\tilde{\Psi}:[0,2\pi]\times[0,2\pi]\rightarrow\mathbb{R}^{4}$ which is defined by
$\tilde{\Psi}(t,s)=\left(f(t),\,(r-d_{c}S_{c}(t))\,g(t)+\cos(s),\,(r-d_{c}S_{c}(t))\,g(t)+\sin(s)+d_{w}S_{w}(t),\,h(t)\right),$
where $r,d_{c},d_{w}\in\mathbb{R}_{+}$ are the radius of the tube, shrinking factor and the displacement factor, respectively, and $h(t)$ provides the over/under information of the tubes in $\mathbb{R}^{4}$. This completes the proof. ∎
Replacing all cosine, sine and bump function with their corresponding Chebyshev approximations, we can get a polynomial parametrization. Mathematica plots of the $Tube(K)$ near the classical and welded crossings are shown in Fig. 21(a) and Fig. 21(a), respectively.
(a) Near classical crossing: lower tube is shrunk
(b) Near welded crossing: upper and lower tubes are displaced
Figure 21. Mathematica plot of $Tube(K)$ near the crossings
4.3. Example: A ribbon torus knot arising from $T(2,7)$
In this example, we provide explicit parametrization of the Tube of the classical torus knot $T(2,7)$ (Fig. 22(a)) with two welded crossings with one gap (Fig. 22(b)). Let’s denote this welded knot as $K_{2,1}$.
(a) Classical $T(2,7)$
(b) Welded knot $K_{2,1}$
Figure 22. Welded knot obtained from $T(2,7)$
The trigonometric parametrization of $T(2,7)$ is given by $\phi^{\prime}(t)=\left(f(t),g(t),h(t)\right)$, where
$f(t)$ $=\cos(2t)(\cos(7t)+3)$
$g(t)$ $=\sin(2t)(\cos(7t)+3)$
$h(t)$ $=\sin(7t).$
The interval pairs $t-\frac{L}{2},t+\frac{L}{2}]$ of equal length $L=\frac{\pi}{7}$ containing the five classical and two welded crossings with over/under information are given in table below.
In this case, we can write
$S_{c}(t)=B\left(t,\frac{11\pi}{14},\frac{\pi}{14}\right)+B\left(t,\frac{15\pi}{14},\frac{\pi}{14}\right)+B\left(t,\frac{19\pi}{14},\frac{\pi}{14}\right)+B\left(t,\frac{23\pi}{14},\frac{\pi}{14}\right)+B\left(t,\frac{27\pi}{14},\frac{\pi}{14}\right),$
$S_{w}(t)=\left(B\left(t,\frac{3\pi}{14},\frac{\pi}{14}\right)+B\left(t,\frac{7\pi}{14},\frac{\pi}{14}\right)\right)-\left(B\left(t,\frac{17\pi}{14},\frac{\pi}{14}\right)+B\left(t,\frac{21\pi}{14},\frac{\pi}{14}\right)\right).$
Choosing $r=0.7,\,d_{c}=1,\,d_{w}=5$, the parametrization of the $\text{Tube}(K_{2,1})$ is given by $\tilde{\Psi}:[0,2\pi]\times[0,2\pi]\rightarrow\mathbb{R}^{4}$ which is defined by
$\tilde{\Psi}(t,s)=\left(f(t),(0.7-S_{c}(t))g(t)+\cos(s),(0.7-S_{c}(t))g(t)+\sin(s)+5S_{w}(t),h(t)\right)$
where $t\in[0,2\pi],s\in[0,2\pi]$. Mathematica plot of the projection of $\text{Tube}(K_{2,1})$ in $XYZ$ hyperspace is shown in Fig. 23.
Figure 23. Projection of $\text{Tube}(K_{2,1})$ in $XYZ$ hyperspace
5. Parametrizing knotted planes
In this section, we develop methods for constructing and parametrizing knotted planes, specifically, long $2$-knots using polynomial embeddings. A long $2$-knot refers to a proper, smooth embedding of $\mathbb{R}^{2}$ into $\mathbb{R}^{4}$, which becomes standard outside a compact region. We begin by introducing polynomial $2$-knots and prove that every such knot can be continuously deformed into a trivial surface knot via a P-homotopy. We then extend the notion of singularity index from long knots to long $2$-knots and establish that every long knot gives rise to a simple long $2$-knot. Finally, we present two geometric constructions of knotted planes via spinning, one through embedded knotted discs and the other by rotating a knotted arc with one end at infinity.
Definition 5.1.
A long $2$-knot is called a polynomial $2$-knot if it is an embedding $\phi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{4}$ defined by
$\phi(t,s)=(f(t,s),g(t,s),h(t,s),p(t,s)),$
where $f(t,s),g(t,s),h(t,s)$ and $p(t,s)$ are real polynomials.
Theorem 5.1.
For every polynomial $2$-knot $F$, there exists a P-homotopy that transforms $F$ to a trivial surface knot.
Proof.
Let $F$ be represented by the the map $\psi:\mathbb{R}^{2}\rightarrow\mathbb{R}^{4}$ defined by
$\phi(t,s)=\big(f(t,s),g(t,s),h(t,s),p(t,s)\big).$
Up to polynomial isotopy we can assume that $p(t,s)$ to be of odd degree. Now, for $u\in\mathbb{R}$, we define a family of maps $\phi_{u}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{4}$ by
$(t,s)\rightarrow\big(f(t,s),g(t,s),h(t,s),p(t,s)+u^{2}(t+s)\big).$
Set $p(t,s)+u^{2}(t+s)=\tilde{p}_{u}(t,s)$. For each $u\in\mathbb{R}$, the Jacobian matrix is
$\begin{pmatrix}\dfrac{\partial f}{\partial t}&\dfrac{\partial f}{\partial s}\\[8.5359pt] \dfrac{\partial g}{\partial t}&\dfrac{\partial g}{\partial s}\\[8.5359pt] \dfrac{\partial h}{\partial t}&\dfrac{\partial h}{\partial s}\\[8.5359pt] \dfrac{\partial p}{\partial t}+u^{2}&\dfrac{\partial p}{\partial s}+u^{2}\end{pmatrix}$
has rank $2$ since $\psi$ is an immersion. Since $p(t,s)$ is of odd degree, all partial derivatives of $p(t,s)$, $\frac{\partial p}{\partial t}$ and $\frac{\partial p}{\partial s}$ are even degree polynomials. As shown in [13], there exist $R_{1},R_{2}\in\mathbb{R}$ such that $\frac{\partial p}{\partial t}+R_{1}>0$ and $\frac{\partial p}{\partial s}+R_{2}>0$
Take $R=\max\{R_{1},R_{2}\}$. Therefore
$\frac{\partial p}{\partial t}+R>0\quad\hfill\text{and}\quad\frac{\partial p}{\partial s}+R>0.$
Let $M=\sqrt{R}$. Therefore, for $u\geq M$,
$\dfrac{\partial\tilde{p}_{u}(t,s)}{\partial t}$ $=\dfrac{\partial p_{u}(t,s)}{\partial t}+u^{2}>0,$
$\text{and}\quad\dfrac{\partial\tilde{p}_{u}(t,s)}{\partial s}$ $=\dfrac{\partial p_{u}(t,s)}{\partial s}+u^{2}>0.$
i.e., $\tilde{p}_{u}(t,s)$is monotonically increasing for $u\geq M$, ensuring that $\phi_{u}$ is a trivial surface knot.
Therefore, we have a homotopy $\phi:\mathbb{R}^{2}\times[0,M]\rightarrow\mathbb{R}^{4}$ defined by
$\phi(t,s,u)=\big(f(t,s),g(t,s),h(t,s),p(t,s)+u^{2}(t+s)\big)$
such that $\phi(t,s,0)=\big(f(t,s),g(t,s),h(t,s),p(t,s)\big)$ is the original long 2-knot $F$ and $\phi(t,s,M)=\big(f(t,s),g(t,s),h(t,s),p(t,s)+M^{2}(t+s)\big)$ presents a trivial long 2-knot. And for each $u\in[0,M]$, $\phi(t,s,u)$ is an immersion. ∎
To understand how such homotopies pass through singular stages, we define the singularity index, generalizing the analogous notion from classical long knots [13].
Remark 5.2.
In Theorem 5.1, for each $u\in[0,M]$ we get an immersion $\phi_{u}$ in $\mathbb{R}^{4}$ with same projection as $F$ in $\mathbb{R}^{3}$. This implies that the singular set is same for each $\phi_{u}$ and the one parameter family $\{p(t,s)+u^{2}(t+s)\,|\,u\in[0,M]\}$ changes the over/under information through the homotopy $\phi$. Therefore, there are finitely many $u\in[0,M]$ for which $\phi_{u}$ is an immersion in $\mathbb{R}^{4}$. Let us refer them as singular long $2$-knots. The number of these singular long $2$-knots in the homotopy $\phi$ is called the singularity index of the polynomial representation $\psi$ of $F$.
Definition 5.2.
Singularity Index of a long $2$-knot $F$ is defined to be the minimum number of singularity indices over all polynomials $\psi$ presenting $F$ and denoted by SI(F).
Definition 5.3.
A long $2$-knot is said to be simple if it has a projection on some $\mathbb{R}^{3}$ such that a singular set consists of finitely many double point curves only.
Next, we establish that any classical polynomial knot can be used to construct a simple long $2$-knot whose singularity index is less than or equal to that of the original knot.
Theorem 5.3.
Given a polynomial knot $K$, with the singularity index $n$ there exists a simple polynomial $2$-knot $P_{K}$ with the singularity index less than or equal to $n$.
Proof.
Let $K:\mathbb{R}\rightarrow\mathbb{R}^{3}$ be a polynomial knot defined as $K(t)=(f(t),g(t),h(t))$ with singularity index $n$, where $(f(t),g(t))$ is an immersion in $\mathbb{R}^{2}$ and $h(t)$ is an odd degree polynomial with positive co-efficient [Remark 2.4, [13]]. Now, we construct a long 2-knot given by $P_{K}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{4}$ and defined by
$P_{K}(t,s)=(f(t)+s,g(t)+s,s,h(t)).$
$K$ being an immersion, the Jacobian matrix for $P_{K}$
$\begin{pmatrix}\dfrac{\partial f}{\partial t}&1\\[8.5359pt] \dfrac{\partial g}{\partial t}&1\\[5.69046pt] 0&1\\[2.84544pt] \dfrac{\partial h}{\partial t}&1\end{pmatrix}$
has rank two. Therefore, $P_{K}$ is an immersion.
Suppose $(t_{1},s_{1})$ and $(t_{2},s_{2})$ be such that
$f(t_{1})+s_{1}$ $=f(t_{2})+s_{2},$
$g(t_{1})+s_{1}$ $=g(t_{2})+s_{2},$
$s_{1}$ $=s_{2},$
$h(t_{1})$ $=h(t_{2}).$
This implies $f(t_{1})=f(t_{2}),\;g(t_{1})=g(t_{2}),\;h(t_{1})=h(t_{2}).$ Since, $K$ is an embedding, we have $t_{1}=t_{2}$. Hence, $(t_{1},s_{1})=(t_{2},s_{2})$ and therefore $P_{K}$ is an injective immersion i.e an embedding. The singular set of the long $2$-knot contains only double curves, one for each double point of $K$. If $(t_{j},t_{k})$ are the values of $t$ for a crossing in the projection of $K$ then the pair $(t_{j},s)$ and $(t_{k},s)$ will provide a double curve in the projection of $P_{K}$ in $\mathbb{R}^{3}$.
By Theorem 5.1, we have a continuous map $\Psi:\mathbb{R}^{2}\times[0,R]\rightarrow\mathbb{R}^{4}$ such that $\Psi(t,s,0)=(f(t)+s,g(t)+s,s,h(t))$ is the given knot $P_{K}$ and $\Psi(t,s,R)=(f(t)+s,g(t)+s,s,h(t)+R^{2}(t+s))$ is a trivial knot and for each $u\in[0,R]$ the map $\Psi((t,s),u)=(f(t)+s,g(t)+s,s,h(t)+u^{2}(t+s))$ is an immersion. Now, restriction of this homotopy $\Psi$ on $\mathbb{R}\times[0,R]$, i.e., $\Psi|_{\mathbb{R}\times[0,R]}\mathbb{R}\times[0,R]\rightarrow\mathbb{R}^{3}$ given by
$(t,u)\rightarrow(f(t)+u,g(t)+u,h(t)+u^{2}t),$
is the homotopy that transforms the long knot $K$ to a trivial classical knot. Now, $K$ has singularity index $n$. Therefore, there are at least $n$ values of $u$, say $u_{1},\cdots u_{n}$ for which $\Psi(t,u_{i})$ is a singular knot. For those values of $u_{i},\,i=1,\cdots,n$,
$f(t_{a})+u_{i}$ $=f(t_{b})+u_{i},$
$g(t_{a})+u_{i}$ $=g(t_{b})+u_{i},$
$h(t_{a})+u_{i}^{2}t_{a}$ $=h(t_{b})+u_{i}^{2}t_{b},$
for all pairs $(t_{a},t_{b})$ representing the crossing points. This implies
$f(t_{a})+u_{i}+s$ $=f(t_{b})+u_{i}+s,$
$g(t_{a})+u_{i}+s$ $=g(t_{b})+u_{i}+s,$
$h(t_{a})+u_{i}^{2}t_{a}$ $=h(t_{b})+u_{i}^{2}t_{b},$
for all the pairs $(t_{a},s)$, $(t_{b},s)$, representing the double curves.
This suggests that $\Psi(t,s,u_{i})$ corresponding to $\Psi(t,u_{i})$ for $u_{i},\,i=1,\cdots,n$ are not embeddings in $\mathbb{R}^{4}$. This indicates the fact that the long $2$-knot $P_{K}$ is transformed to the trivial surface knot diagram by exchanging over/under information in the set of $n$ number of double curves. Hence, the singularity index of $P_{K}$ is less or equal to $n$. ∎
Example 5.1.
Here is an example of a long $2$-knot (Fig. 24) constructed from the long trefoil given by the map
$(t,s)\to(t^{3}-3t+s,t^{4}-4t^{2}+s,s,t^{5}-10t).$
Figure 24. Projection of the simple long trefoil 2-knot on a hyperplane
We now present two constructions of knotted planes in $\mathbb{R}^{4}$, both derived from a given long knot $K$ using spinning operations.
5.1. Constructing knotted planes by spinning
Construction I: This construction starts with a knotted disc bounded by the connected sum $K\sharp K^{*}$ and extends it to a full knotted plane via a polynomial homotopy.
Lemma 5.4.
For every polynomial $K$, there exists a knotted disc bounded by $K\sharp K^{*}$.
Proof.
First, take a knotted arc $k$ of the polynomial knot $K$ in the upper half space $\mathbb{R}^{3}_{+}$ with only two endpoints on the boundary $\mathbb{R}^{2}$. Let the knotted arc be presented by $\big(f(t),g(t),h(t)\big)$ where $t\in[a,b]$ where $h(a)=0=h(b)$ and $h(t)>0\quad\text{for}\,t\in(a,b)$. Spinning this arc around $\mathbb{R}^{2}$ from angle 0 to $2\pi$, results into a semi sphere or a $2$-disc in $\mathbb{R}^{4}$ bounded by $K\sharp K^{*}$ where $K^{*}$ is the mirror image of $K$. Denote it by $D$. Then $D$ can be parametrized by
$a,b]\times[0,\pi]\rightarrow\mathbb{R}^{4}$
$(t,s)\rightarrow\left(f(t),g(t),h(t)\sin s,h(t)\cos s\right),$
where $s=0$ and $s=\pi$ will map to the knotted arcs $k$ and $k^{*}$ on the boundary of $D$, given by
$\left(f(t),g(t),0,h(t)\right)\,\text{for}\,t\in[a,b],$
$\left(f(t),g(t),0,-h(t)\right)\,\text{for}\,t\in[a,b]$
respectively. ∎
Example 5.2.
Here is an example of a knotted disc with boundary $K\sharp K^{*}$ where $K$ is the long trefoil (Fig. 25), given by
$a,b]\times[0,\pi]\rightarrow\mathbb{R}^{4}$
$(t,s)\rightarrow\left(t^{3}-3t,t^{5}-10t,(-t^{4}+4t^{2}+3)\sin s,(-t^{4}+4t^{2})+3)\cos s\right)$
Figure 25. Knotted disc with boundary $K\sharp K^{*}$
Theorem 5.5.
For every polynomial knot $K$, there exists a knotted plane that contains an knotted disc bounded by $K\sharp K^{*}$.
Proof.
By Lemma 5.4, we have a knotted disc $D$ with boundary $K\sharp K^{*}$, given by $\phi:[a,b]\times[0,2\pi]\rightarrow\mathbb{R}^{4},$ where
$\phi(t,s)=\big(f(t),g(t),h(t)\sin s,h(t)\cos s\big).$
Now we construct an annulus $A$ with boundary $K\sharp K^{*}$ in one end and an unknotted $S^{1}$ in the other end by using the P-homotopy map mentioned in Theorem 5.3. For $t\in[a,b],\,s\in[-\infty,0]$, we define
$P_{K}:$ $\;\psi_{1}(t,s)=\big(f(t)-s,g(t)-s,s,h(t)+s^{2}t\big),$
$P_{K^{*}}:$ $\;\psi_{2}(t,s)=\big(f(t)+(s-\pi),g(t)+(s-\pi),-(s-\pi),-h(t)+(s-\pi)^{2}t\big),$
where $\psi_{1}(t,s)$ and $\psi_{2}(t,s)$ are trivial diagrams for all $s\leq-R$ and $s\geq\pi+R$ by [Lemma 2.7, [13]]. Also, we have the following:
  • $\psi_{1}(t,0)=\phi(t,0)$, $\psi_{2}(t,\pi)=\phi(t,\pi)$ presents $k$ and $k^{*}$ respectively.
  • $k$ and $k^{*}$ are joined at the endpoints $\psi_{1}(a,0)=\psi_{1}(a,0)$, $\psi_{2}(b,\pi)=\psi_{2}(t,\pi)$.
  • The values of $\psi_{1}(a,s)$ for $s\in[-\infty,0]$ and $\psi_{2}(b,s)$ for $s\in[\pi,\infty]$ are equal since $h(a)=0=h(b)$. Therefore, $P_{K}$ and $P_{K^{*}}$ are joined along the lines $\psi_{1}(a,s),\,s\in[-\infty,0]$ and $\psi_{2}(b,s),\,s\in[\pi,\infty]$.
  • Then the annulus $A$ is obtained by restricting $P_{K}\,\cup\,P_{K^{*}}$ to $s\in[-R,0]\cup[0,R]$.
  • The other end of the annulus $A$ is a unknotted circle created by $\psi_{1}(t,-R)$ and $\psi_{2}(t,\pi+R)$ which are unknotted arcs joined at the endpoints $\psi_{1}(a,-R)=\psi_{1}(a,\pi+R)$ and $\psi_{2}(a,-R)=\psi_{2}(a,\pi+R)$.
  • Outside the interval $s\in[-R,R]$, $\psi_{1}(t,s)\cup\psi_{2}(t,s)$ will always be a unknotted $S^{1}$.
So, $D\cup A$ joined along the boundary $K\sharp K^{*}$ gives us an embedding of an open knotted disc with boundary an unknotted $S^{1}$. Therefore, extending the interval for $s$ from [R,R][-R,R] to $\mathbb{R}$ we get a parametrization of a knotted plane given by
$\Phi:(a,b)\times\mathbb{R}\rightarrow\mathbb{R}^{4}$
$\Phi=\begin{cases}\psi_{1}&s\in(-\infty,0]\\ \phi&s\in[0,\pi]\\ \psi_{2}&s\in[\pi,\infty)\end{cases}$
This completes the proof. ∎
Remark 5.6.
This construction shows that $K\sharp K^{*}$ bounds a smooth disc in $D^{4}$, hence it is a slice knot. [4].
Example 5.3.
For the long trefoil knot $K$, we have the knotted plane given by $\Phi:(a,b)\times\mathbb{R}\rightarrow\mathbb{R}^{4},$ which is defined by
$\Phi=\begin{cases}\psi_{1}&s\in(-\infty,0],\\ \phi&s\in[0,\pi],\\ \psi_{2}&s\in[\pi,\infty),\end{cases}$
$\text{where}\quad\phi(t,s)$ $=\big(t^{3}-3t,t^{5}-10t,(-t^{3}+4t^{2}+3)\sin s,(-t^{3}+4t^{2}+3)\cos s\big),$
$\psi_{1}(t,s)$ $=\big(t^{3}-3t-s,t^{5}-10t-s,s,(-t^{3}+4t^{2}+3)+s^{2}t\big),$
$\psi_{2}(t,s)$ $=\big(t^{3}-3t+(s-\pi),t^{5}-10t+(s-\pi),-(s-\pi),(t^{3}-4t^{2}-3)+(s-\pi)^{2}t\big).$
In Fig. 26, $D$ is shown in orange color and the annulus $A$ is shown in blue and green for $P_{k}$ and $P_{K^{*}}$ respectively.
Construction II: This method spins an arc of a long knot $K$ extending its one end to the infinity in $\mathbb{R}^{3}_{+}$ to produce a knotted plane in $\mathbb{R}^{4}$.
Figure 26. Knotted plane with a knotted disc bounded by $K\sharp K^{*}$.
Theorem 5.7.
For a polynomial knot $K$ given by $\big\{(f(t),g(t),h(t))\,|\,t\in\mathbb{R}\big\}$, there exists a polynomial 2-knot obtained from spinning operation with parametrization given by $\psi^{\prime}:[a,\infty]\times[0,2\pi]\rightarrow\mathbb{R}^{4}$ which is defined by
$\psi^{\prime}(t,s)=\left(f_{1}(t),g_{1}(t),h_{1}(t)\,S(s),h_{1}(t)\,C(s)\right).$
Proof.
We start with the knotted arc of $K$ in the upper half space $\mathbb{R}^{3}_{+}$. Instead of taking both endpoints of the knotted arc on $\mathbb{R}^{2}$, we put one endpoint on $\mathbb{R}^{2}$, and the other end goes to infinity. This happens when $h(t)$ is an odd degree polynomial. If $h(t)$ is an even degree, then we can add a term of a larger odd degree with a very small coefficient so that the over/under information is not disturbed (Fig. 27). Let this knotted arc be given by
$a,\infty)\rightarrow\mathbb{R}^{3}$
$t\rightarrow\left(f_{1}(t),g_{1}(t),h_{1}(t)\right).$
Figure 27. Long trefoil arc for $t\in[a,\infty]$
Now, rotating this knotted arc around $\mathbb{R}^{2}$ results in a knotted plane in $S^{4}$ which can be parametrized by
$a,\infty)\times[0,2\pi]\rightarrow\mathbb{R}^{4}$
$t\rightarrow\big(f_{1}(t),g_{1}(t),h_{1}(t)\sin s,h_{1}(t)\cos s\big).$
Replacing $\cos s$ and $\sin s$ with their corresponding Chebyshev approximations $C(s)$ and $S(s)$, respectively, we get the final polynomial parametrization. This completes the proof. ∎
Example 5.4.
An example of a knotted plane obtained from a long trefoil is shown in Fig. 28.
(a) The knotted part inside a compact region
(b) Inside view
Figure 28. Knotted plane obtained by spinning
6. Conclusion:
Our constructions provide infinitely many example of knotted surfaces. In case of knotted spheres and knotted tori these surface knots are compact. However, Knotted planes are non-compact. We have seen that their constructions rely on classical knots, once the parametrization of classical knots is known these knotted surfaces are easily parametrized and their geometric properties can be studied. One future goal is to extend parametrization to broader classes of knotted surfaces. This includes non‐orientable embeddings (like projective planes or Klein bottles in four-dimensional space), surfaces of higher genus, or surfaces with branch points. Moreover, a degree‐minimization problem i.e finding the lowest polynomial degree that realizes a given knotted surface is another problem to study in future.
7. Acknowledgment
The second author gratefully acknowledges financial support from the Council of Scientific and Industrial Research (CSIR), India.